OpenStudy (anonymous):

Find all points on the curve r(t)= (t, t^2,t^3) where its tangent line is parallel to the vector v = i +4j +12k . Anyone help me, pleae

OpenStudy (jamesj):

So if the tangent to that curve is parallel to the vector v, what can you say? Call the tangent T. T is parallel to v is there is a constant k such that T = kv, yes? Now, T changes with time T(t). You find it by differentiating r(t); i.e., T(t) = r'(t) Now see what that implies for different values of t.

OpenStudy (anonymous):

yes. I got \[T(t)=\frac{ <1,2t,3t^2> }{ \sqrt{1+4t^2+9t^4}? }\]

OpenStudy (jamesj):

No need to normalize the T vector

OpenStudy (anonymous):

and then break it down under the form of a vector. since vector v = <1,4,12> I let \[\frac{ 1 }{\sqrt{1+4t^2+9t^4} } = k*1(1 from x value of vector v)\]

OpenStudy (jamesj):

yes. But again, loose the normalization, as it isn't necessary and it will significantly simplify your calculations.

OpenStudy (anonymous):

and let Ty =k Vy Tz =kVz and solve the system. if you don't break the T, how to find out the k

OpenStudy (anonymous):

please show me the stuff. I really stuck there

OpenStudy (jamesj):

using the unnormalized tangent vector <1,2t,3t^2> = k<1,4,12> Hence 1 = k 2t = 4k 3t^2 = 12k Now, for what values of t can this be true?

OpenStudy (anonymous):

2

OpenStudy (jamesj):

t = 2, yes. Hence there is only one point on the curve where the tangent vector is parallel to v, and that is r(t=2) = .....

OpenStudy (anonymous):

can we ignore the denominator of T(t)? because that vector totally different from r(t)' which is just numerator of T(t)

OpenStudy (jamesj):

All we are checking is whether two vectors are parallel. In general, if v and w are parallel, then av and bw are also parallel for any non-zero constants a and b. So here, we are simplifying the calculations by just using the unnormalized form of T.

OpenStudy (jamesj):

in fact, v and w are parallel if and only if av and bw are parallel for any non-zero constants a and b.

OpenStudy (anonymous):

Thank you very much. can you give me some thing related to "unnormalized" form as conference to read (some thing like theorem, article or else.) i want to make sure 100% understand what i am doing. anyway, thanks a lot